Find a formula that generates the following sequence a1,a2,a3...
-1, 1, 7, 17, 31, 49, 71
Answer #1
f(1) = -1
f(n) = 2 + f(n-1) + 4*(n-2)
     = 2 + f(n-1) + 4*(n-2)
     = 2 + f(n-1) + 4n-8
     = f(n-1) + 4n - 6

f(1) = -1
f(n) = f(n-1) + 4n - 6